One angle of a triangle measures 110 degrees. The other two angles are in
a ratio of 2:5. What are the measures of the other two angles?

Answer by LinnW(1048)

The remaining two angles total (180-100) = 70 degrees.
With a ratio of 2:5 , the "whole" is the sum of 2 and 5 = 7
So the first angle should be 2/7 and the second should be 5/7
of the 70 degrees.
2/7 * 70 = 20 degrees
5/7 * 70 = 50 degrees